### v2

# LLMs cannot find reasoning errors, but can correct them!

Abstract
While self-correction has shown promise in improving LLM outputs in terms of style and quality (e.g. Chen et al., 2023; Madaan et al., 2023), recent attempts to self-correct logical or reasoning errors often cause correct answers to become incorrect, resulting in worse performances overall (Huang et al., 2023). In this paper, we break down the self-correction process into two core components: mistake finding and output correction. For mistake finding, we release BIG-Bench Mistake, a dataset of logical mistakes in Chain-of-Thought reasoning traces. We provide benchmark numbers for several state-of-the-art LLMs, and demonstrate that LLMs generally struggle with finding logical mistakes. For output correction, we propose a backtracking method which provides large improvements when given information on mistake location. We construe backtracking as a lightweight alternative to reinforcement learning methods, and show that it remains effective with a reward model at 60-70% accuracy.
1 Introduction
Large Language Models (LLMs) have dominated the field of NLP in recent years, achieving state-of-the-art performance in a large variety of applications. In particular, LLMs have demonstrated the ability to solve tasks with zero- or few-shot prompting, giving rise to prompting methods such as Chain-of-Thought (CoT) (Wei et al., 2022), Self-Consistency (SC) (Wang et al., 2023), ReAct (Yao et al., 2022), etc.
Recent literature on few- or zero-shot prompting has focused on the concept of self-correction, i.e. having an LLM correct its own outputs (Shinn et al., 2023; Miao et al., 2023; Madaan et al., 2023; Chen et al., 2023; Saunders et al., 2022). (See Pan et al. (2023) for a review of the literature.)
However, Huang et al. (2023) note that while self-correction may prove effective for improving model outputs in terms of style and quality, there is limited evidence that LLMs can identify and fix their own reasoning and logical errors without external feedback. For example, Reflexion (Shinn et al., 2023) and RCI (Kim et al., 2023) both use ground truth correctness as a signal to halt the self-correction loop. Initially observed by Madaan et al. (2023) on a math reasoning dataset, Huang et al. (2023) further demonstrate this shortcoming of self-correction in 2 additional datasets.
While previous work typically present self-correction as a single process, we divide it into mistake finding and output correction.
Mistake finding is a fundamental reasoning skill that has been studied and utilised extensively in philosophy, psychology, and mathematics, spawning concepts such as critical thinking, and logical and mathematical fallacies. One might expect that the ability to find mistakes should also be an important requirement for LLMs. However, our results show that state-of-the-art LLMs currently cannot find mistakes reliably.
Output correction involves partially or completely changing previously generated outputs. In the context of self-correction, this is typically done with outputs generated by the same model (see Pan et al. (2023) for an overview of different strategies). Despite LLMs’ inability to find mistakes, our results show that they can correct outputs using our backtracking method, if given information about the mistakes, for example via a small, supervised reward model.
Our contributions for this paper are as follows:
1.
With Chain-of-Thought prompting, any task can be turned into a mistake-finding task. We collect and release111https://github.com/WHGTyen/BIG-Bench-Mistake. to the research community BIG-Bench Mistake, a dataset of CoT-style traces222In this paper, we refer to a set of CoT reasoning steps as a trace. generated using PaLM 2, and annotated according to where the first logical mistake is. To our knowledge, BIG-Bench Mistake is the first dataset of its kind that goes beyond problems in mathematics.
2.
We produce benchmark results for our dataset to test the reasoning capabilities of state-of-the-art LLMs. We demonstrate that current state-of-the-art LLMs struggle with mistake finding, even for objective, unambiguous cases. We hypothesise that this is a main contributing factor to LLMs’ inability to self-correct reasoning errors, and we call on the research community to pursue further improvements on the mistake finding task.
3.
We propose backtracking as an output correction technique that makes use of mistake location information to improve performance on the original task. We demonstrate that this method corrects outputs that are originally incorrect, with minimal effect on outputs that are originally correct.
4.
We construe backtracking as a form of “verbal reinforcement learning” (Shinn et al., 2023), allowing iterative improvement on CoT outputs without requiring any weight updates. We propose that backtracking can be used with a trained classifier as a reward model, and demonstrate the effectiveness of backtracking at various reward model accuracies.
2 BIG-Bench Mistake
BIG-Bench Mistake consists of 2186 sets of CoT-style traces. Each trace was generated by PaLM 2-L-Unicorn, and annotated with the location of the first logical error. An example trace is shown in Table 1, where the mistake location333As some traces may not contain mistakes, we use the term mistake location as a multi-class label that can refer to either the integer where the th step contains the first mistake, or that there are no mistakes. is the 4th step.
| Sort the following words alphabetically: List: hypochlorite ponderosa phone credulity | |
| Thought 1: |
I should start by looking at the first letter of the words in the list. The first letter: "hypochlorite": "h" (8). "ponderosa": "p" (16). "phone": "p" (16). "credulity": "c" (3). |
| Thought 2: |
We now have: (3) "credulity" < (8) "hypochlorite" < (16) ["ponderosa" ? "phone"]. |
| Thought 3: |
Now let’s sort this subpart ["ponderosa" ? "phone"] by looking at their second letters. The second letter: "ponderosa": "o" (15). "phone": "h" (8). |
| Thought 4: |
We now have: (8) "phone" < (15) "ponderosa" for the subpart. Hence, we have "credulity" < "phone" < "ponderosa". (MISTAKE) |
| Thought 5: |
I have now sorted all the words. The answer is credulity hypochlorite phone ponderosa |
Our traces span across a set of 5 tasks444These 5 tasks were selected because 1) Anil et al. (2023) demonstrate that PaLM 2 performs poorly on these tasks, so it is likely to generate mistakes in CoT traces; 2) any mistakes that may occur are likely to be unambiguous, therefore minimising subjectivity during annotation; and 3) identifying mistakes for these tasks does not require expertise knowledge of a specific domain. from the BIG-bench dataset (Srivastava et al., 2023): word sorting, tracking shuffled objects, logical deduction, multi-step arithmetic, and Dyck languages. CoT prompting is used to prompt PaLM 2 to answer questions from each task. As we wanted to separate our CoT traces into distinct steps, we follow the method used by Yao et al. (2022) and generate each step separately, using the newline as a stop token.
In this dataset, all traces are generated with temperature . The correctness of answers are determined by exact match. Prompts can be found at https://github.com/WHGTyen/BIG-Bench-Mistake.
2.1 Annotation
Each generated trace is annotated with the first logical error. We ignore any subsequent errors as they may be dependent on the original error.
Note that traces can contain a logical mistake yet arrive at the correct answer. To disambiguate the two types of correctness, we will use the terms correct and incorrect to refer to whether the final answer of the trace is correct. Accuracy would therefore refer to the overall accuracy for the task, based on how many final answers are correct. To refer to whether the trace contains a logical mistake (rather than the correctness of the final answer), we will use correct and incorrect.
2.1.1 Human annotation
For 4 of the 5 tasks, we recruit human annotators to go through each trace and identify any errors. Annotators have no domain expertise but are given guidelines555See https://github.com/WHGTyen/BIG-Bench-Mistake for further details. to complete the task.
Before annotation, we sample a set of 300 traces for each task, where 255 (85%) are incorrect, and 45 (15%) are correct. Since human annotation is a limited and expensive resource, we chose this distribution to maximise the number of steps containing mistakes and to prevent over-saturation of correct steps. We also include some correct traces because some may contain logical errors despite the correct answer, and to ensure that the dataset included examples of correct steps that are near the end of the trace. This also prevents annotators from feeling forced to find a mistake in all traces.To account for this skewed distribution, results in section 4 are split according to whether the original trace is correct or incorrect.
Following Lightman et al. (2023), annotators are instructed to go through each step in the trace and indicate whether the step is correct or not (binary choice). Annotators only submit their answers until all steps have been annotated, or there is one incorrect step. If an incorrect step is identified, the remaining steps are skipped. This is done to avoid ambiguities where a logically correct deduction is dependent on a previous mistake. We make our annotation guidelines available at https://github.com/WHGTyen/BIG-Bench-Mistake, and we include a screenshot of the user interface in Figure 3.
Each trace has been annotated by at least 3 annotators. If there are any disagreements, we take the majority label. We calculate Krippendorff’s alpha (Hayes and Krippendorff, 2007) to measure inter-rater reliability (see Table 2).
| Task | Krippendorff’s |
| Word sorting | 0.979 |
| Tracking shuffled objects | 0.998 |
| Logical deduction | 0.996 |
| Multistep arithmetic | 0.984 |
2.1.2 Automatic annotation
For Dyck languages, we opt for mostly automatic annotation instead of human annotation as the traces show limited variation in phrasing and solution paths.
For each trace, we generate a set of standard steps based on the format used in the prompt examples. Using pattern matching, we can identify whether each model-generated step also conforms to the same format. If so, we compare the two and assume that the trace is incorrect if the symbols do not match. Additionally, we also account for edge cases such as where the final two steps are merged into one, or variations in presentation where some symbols are placed in quotes and some are not. We release the code at https://github.com/WHGTyen/BIG-Bench-Mistake along with our dataset.
| Task |
Num. of correct traces |
Num. of incorrect |
Num. of incorrect traces |
Total |
| Word sorting |
45 |
255 |
266 |
300 |
| Tracking shuffled objects |
45 |
255 |
260 |
300 |
| Logical deduction |
45 |
255 |
294 |
300 |
| Multistep arithmetic |
45 |
255 |
238 |
300 |
| Dyck languages |
482 |
504 |
650 |
986 |
| Dyck languages (sampled) |
88 |
504 |
545 |
592 |
3 Benchmark results
Table 4 shows the accuracy of GPT-4-Turbo, GPT-4, and GPT-3.5-Turbo on our mistake-finding dataset. For each question, the possible answers are either that there are no mistakes, or, if there is a mistake, the number N indicating the step in which the first mistake occurs. A model’s output is only considered correct if the location matches exactly, or the output correctly indicates that there are no mistakes.
All models are given the same 3-shot prompts666Prompts can be found at https://github.com/WHGTyen/BIG-Bench-Mistake.. We use three different prompting methods:
Direct trace-level prompting involves using the whole trace as input to the model and directly prompting for the mistake location. The model must output either the number representing the step, or "No".
Direct step-level prompting prompts for a binary Yes/No output for every step, indicating whether or not the step is correct. In each generation call, the input contains the partial trace up to (and including) the target step, but does not contain results for previous steps. The final answer is inferred from where the first "No" output occurs (subsequent steps are ignored).
CoT step-level prompting is an extension of direct, step-level prompting. Instead of a binary Yes/No response, we prompt the model to check the (partial) trace through a series of reasoning steps. This method is the most resource intensive of all three methods as it involves generating a whole CoT sequence for every step. Due to cost and usage limits, we are unable to provide results from GPT-4-Turbo here. As with direct step-level prompting, the final answer is inferred from where the first "No" output occurs (subsequent steps are ignored).
3.1 Discussion
All three models appear to struggle with our mistake finding dataset. GPT-4 attains the best results but only reaches an overall accuracy of 52.87 with direct step-level prompting.
Our findings are in line with and builds upon results from Huang et al. (2023), who show that existing self-correction strategies are ineffective on reasoning errors. In our experiments, we specifically target the models’ mistake finding ability and provide results for additional tasks. We show that state-of-the-art LLMs clearly struggle with mistake finding, even in the most simple and unambiguous cases. (For comparison, humans can identify mistakes without specific expertise, and have a high degree of agreement, as shown in Table 2.)
We hypothesise that LLMs’ inability to find mistakes is a main contributing factor to why LLMs are unable to self-correct reasoning errors. If LLMs are unable to identify mistakes, it should be no surprise that they are unable to self-correct either.
Note that the mistakes in our dataset are generated using PaLM 2 L (Unicorn), and traces were sampled according to whether the final answer was correct or not. Therefore, we expect that using PaLM 2 itself to do mistake finding will produce different and likely biased results. Further work is needed to elucidate the difference between cross-model evaluation and self-evaluation.
| Model |
Direct (trace) |
Direct (step) |
CoT (step) |
| Word sorting (11.7) | |||
| GPT-4-Turbo |
36.33 |
33.00 |
| GPT-4 |
35.00 |
44.33 |
34.00 |
| GPT-3.5-Turbo |
11.33 |
15.00 |
15.67 |
| Tracking shuffled objects (5.4) | |||
| GPT-4-Turbo |
39.33 |
61.67 |
| GPT-4 |
62.29 |
65.33 |
90.67 |
| GPT-3.5-Turbo |
10.10 |
1.67 |
19.00 |
| Logical deduction (8.3) | |||
| GPT-4-Turbo |
21.33 |
75.00 |
| GPT-4 |
40.67 |
67.67 |
10.33 |
| GPT-3.5-Turbo |
2.00 |
25.33 |
9.67 |
| Multistep arithmetic (5.0) | |||
| GPT-4-Turbo |
38.33 |
43.33 |
| GPT-4 |
44.00 |
42.67 |
41.00 |
| GPT-3.5-Turbo |
20.00 |
26.00 |
25.33 |
| Dyck languages† (24.5) | |||
| GPT-4-Turbo |
15.33* |
28.67* |
| GPT-4 |
17.06 |
44.33* |
41.00* |
| GPT-3.5-Turbo |
8.78 |
5.91 |
1.86 |
| Overall | |||
| GPT-4-Turbo |
30.13 |
48.33 |
| GPT-4 |
39.80 |
52.87 |
43.40 |
| GPT-3.5-Turbo |
10.44 |
14.78 |
14.31 |
† indicates that traces were sampled to contain 15% correct and 85% incorrect traces (see Table 3).
* indicates that traces were sampled to contain 45 correct and 255 incorrect traces to reduce costs.
3.2 Comparison of prompting methods
As we compare results across the three methods, we find that the accuracy on traces with no mistakes goes down777Note that the traces in BIG-Bench Mistake are sampled to contain more incorrect traces than correct traces (and therefore more incorrect traces than correct traces), so the overall mistake location accuracy appears higher for per-step prompting in Table 4, despite the poor accuracy for correct traces. For a full set of split by correctness, see Appendix B. considerably from direct, trace-level prompting to CoT, step-level prompting. Figure 1 demonstrates this trade-off.
We hypothesise that this is due to the number of outputs generated by the model. Our three methods involve generating increasingly complex outputs, starting with direct, trace-level prompting requiring a single token, then direct, step-level prompting requiring one token per step, and finally CoT step-level prompting requiring several sentences per step. If each generation call has some probability of identifying a mistake, then the more calls made on each trace, the more likely the model will identify at least one mistake.
3.3 Few-shot prompting for mistake location as a proxy for correctness
In this section, we investigate whether our prompting methods can reliably determine the correctness of a trace rather than the mistake location. Our motivation was that even humans use mistake finding as a strategy for determining whether an answer is correct or not, such as when going through mathematical proofs, or working through argumentation. Additionally, one might think that directly predicting the correctness of a trace may be easier than having to pinpoint the precise location of an error.
We calculate averaged F1 scores based on whether the model predicts that there is a mistake in the trace. If there is a mistake, we assume the model prediction is that the trace is incorrect. Otherwise, we assume the model prediction is that the trace is correct.
| Model |
Direct (trace) |
Direct (step) |
CoT (step) |
| Word sorting | |||
| GPT-4-Turbo |
87.73 |
86.68 |
| GPT-4 |
81.50 |
85.12 |
81.19 |
| GPT-3.5-Turbo |
6.58 |
35.07 |
77.79 |
| Tracking shuffled objects | |||
| GPT-4-Turbo |
52.23 |
74.31 |
| GPT-4 |
76.38 |
75.69 |
95.03 |
| GPT-3.5-Turbo |
32.04 |
77.61 |
78.11 |
| Logical deduction | |||
| GPT-4-Turbo |
86.46 |
81.79 |
| GPT-4 |
84.54 |
83.38 |
23.96 |
| GPT-3.5-Turbo |
10.34 |
67.62 |
61.31 |
| Multistep arithmetic | |||
| GPT-4-Turbo |
71.17 |
86.24 |
| GPT-4 |
72.97 |
78.67 |
79.67 |
| GPT-3.5-Turbo |
3.76 |
53.18 |
64.08 |
| Dyck languages | |||
| GPT-4-Turbo |
51.96 |
85.87 |
| GPT-4 |
62.33 |
85.73 |
79.60 |
| GPT-3.5-Turbo |
46.57 |
79.31 |
77.79 |
In Table 5, we average the F1s using correct and incorrect as the positive label, weighted according to the number of times each label occurs. Note that the baseline of predicting all traces as incorrect achieves a weighted F1 average of 78.
The weighted F1 scores show that prompting for mistakes is a poor strategy for determining the correctness of the final answer. This is in line with our previous finding that LLMs struggle to identify mistake locations, and also builds upon results from Huang et al. (2023), who demonstrate that improvements from Reflexion (Shinn et al., 2023) and RCI (Kim et al., 2023) are only from using oracle correctness information.
4 Backtracking
Madaan et al. (2023) and Huang et al. (2023) both demonstrate that self-correction is only effective with external feedback – for example, both Shinn et al. (2023) and Kim et al. (2023) rely on oracle labels for improvements – but there is often no external feedback available in many real-world applications.
As an alternative, we explore the possibility of replacing external feedback with a lightweight classifier trained on a small amount of data. Analogous to reward models in conventional reinforcement learning, this classifier detects any logical errors in a CoT trace, which is then fed back to the generator model to improve on the output. This can be done over multiple iterations to maximise improvements.
We propose a simple backtracking method to improve model outputs based on the location of logical errors:
1.
First, the model generates an initial CoT trace. In our experiments, we use temperature .
2.
We then determine the mistake location in this trace using a reward model.
3.
If there are no mistakes, we move onto the next trace. If there is a mistake (e.g. at Thought 4 in the example trace in Table 1),we prompt the model again for the same step but at temperature , generating 8 outputs. We use same prompt and the partial trace containing all steps up to but not including the mistake step (e.g. up to Thought 3, prompting for Thought 4).
4.
From the 8 outputs, we filter out any options that match what was previously identified as a mistake. From the remaining outputs, we select one with the highest log-probability.
5.
Finally, with the new, regenerated step in place of the previous one, we generate the remaining steps of the trace again at temperature .
Our backtracking method provides several benefits over existing self-correction methods:
Unlike Shinn et al. (2023), Kim et al. (2023), etc., our approach does not depend on oracle knowledge of the answer. Instead, it relies on information (for example from trained a reward model) about logical errors, which can be determined on a step-by-step basis using a reward model. Logical errors can occur in correct traces, or not occur in incorrect traces888Having no logical errors in incorrect traces is much rarer but does exist, for example when the answer is correct but is not captured by exact match, or if the original question is faulty and has multiple possible answers..
Compared to approaches that require re-generating the entire trace, backtracking reduces computational cost by reusing previous steps that are known to be logically sound.
Backtracking improves on the quality of the intermediate steps directly, which can be useful in scenarios that require correct steps (e.g. generating solutions to math questions), and also generally improves interpretability.
Backtracking with mistake location information from a reward model can be construed as a lightweight RL method. However, unlike conventional deep reinforcement learning:
Backtracking with a reward model does not does not require any training of the original generator model. Once the reward model is trained, it can be used for backtracking with any LLM as the generator, and can also be updated independently of the generator LM. This can be especially helpful when LLMs are frequently updated to new checkpoints.
Backtracking only requires training of a small reward model. Compared to methods that require training of the generator model, backtracking is far more efficient in terms of computing resources and available data.
The process of backtracking is more interpretable than updating the weights of the generator model directly, as is required for many deep RL methods. It clearly pinpoints the location at which an error occurs, which can help the debugging process and allow faster development and iterations of models.
4.1 Backtracking with gold mistake location
As an initial experiment, we use labels from BIG-Bench Mistake to test if an LLM is able to correct logical errors using backtracking, independent of its inherent ability to identify these errors or any other reward model.
For example, if the mistake location is in step 4, we use backtracking to regenerate that step and continue the rest of the chain. If the mistake location is that there are no logical mistakes, we do not backtrack and use the original result.
| With mistake location | With random location |
Avg. num. of steps |
| Task | accuracy | accuracy | accuracy | accuracy | |
| Word sorting | -11.11 | +23.53 | -15.56 | +11.76 |
11.7 |
| Tracking shuffled objects | -6.67 | +43.92 | -6.67 | +20.39 |
5.4 |
| Logical deduction | -11.43 | +36.86 | -13.33 | +21.57 |
8.3 |
| Multistep arithmetic | -0.00 | +18.04 | -8.89 | +10.59 |
5.0 |
| Dyck languages | -6.82 | +18.06 | -15.91 | +5.16 |
24.5 |
4.1.1 Results
The results are shown in Table 6. To show that performance increases are not due to randomly resampling outputs, we compare our results to a random baseline, where a mistake location999As described above, the mistake location can be either the number representing the step, or that there are no mistakes. If there are no mistakes, we do not use backtracking and simply use the original trace. is randomly selected for each trace and we perform backtracking based on the random location.
Note that Table 6 separates results into numbers for the correct set and the incorrect set, referring to whether the original trace was correct or not. This gives a clearer picture than the overall accuracy, which would be skewed by the proportion of traces that were originally correct (15%) and incorrect (85%).
Scores represent the absolute differences in accuracy. We perform backtracking on both correct and incorrect traces, as long as there is a mistake in one of the steps.
accuracy
refers to differences in accuracy on the set of traces whose original answer was correct. Note that we take losses here because, despite the correct answer, there is a logical mistake in one of the steps. Therefore, the answer may change to an incorrect one when we backtrack.
accuracy
is the same but for incorrect traces, so the answers may have been corrected, hence increasing accuracy.
For example, for the word sorting task, 11.11% of traces that were originally correct became incorrect, while 23.53% of traces that were originally incorrect became correct.
4.1.2 Discussion
The scores show that the gains from correcting incorrect traces are larger than losses from changing originally correct answers. Additionally, while the random baseline also obtained improvements, they are considerably smaller than if the true mistake location was used. Note that tasks involving fewer steps are more likely to improve performance in the random baseline, as the true mistake location is more likely to be identified.
While our numbers do show that our gains are higher than our losses, it should be noted that changes in the overall accuracy depends on the original accuracy achieved on the task. For example, if the original accuracy on the tracking shuffled objects task was 50%, the new accuracy would be 68.6%. On the other hand, if the accuracy was 99%, the new accuracy would drop to 92.8%. As our dataset is highly skewed and only contains 45 correct traces per task, we leave to future work to assess the effectiveness of backtracking in a more comprehensive way.
4.2 Backtracking with a simulated reward model
We show in subsection 4.1 that backtracking can be used to correct CoT traces using gold mistake location labels. To explore what level of accuracy reward model is needed when gold labels are not available, we use backtracking with simulated reward models, designed to produce labels at different levels of accuracy. We use accuracy to refer to the accuracy of the simulated reward model at identifying mistake locations.
For a given reward model at % accuracy, we use the mistake location from BIG-Bench Mistake % of the time. For the remaining %, we sample a mistake location randomly. To mimic the behaviour of a typical classifier, mistake locations are sampled to match the distribution found in the dataset. We also ensure that the sampled location does not match the correct location.
4.2.1 Results
Results are shown in Figure 2. We can see that the losses in accuracy begins to plateau at 65%. In fact, for most tasks, accuracy is already larger than accuracy at around 60-70% accuracy. This demonstrates that while higher accuracies produce better results, backtracking is still effective even without gold standard mistake location labels.
4.3 Reward Modeling
We perform a preliminary investigation of if mistake-finding can benefit from a dedicated reward model and if learning to find mistakes in a set of tasks can transfer to finding mistakes in out-of-distribution tasks. We fine-tuned a PaLM 2-XS-Otter model based on our available data for 20k steps and choose the checkpoint with the best validation results. We hold out one task for evaluation while training the reward model on the other 4 tasks.
| Held-out task | Difference |
| Word sorting | +11.66 |
| Tracking shuffled objects | +19.67 |
| Logical deduction | -0.67 |
| Multi-step arithmetic | +4.00 |
| Dyck languages | +22.59 |
Note the reward model we train is significantly smaller than our inference model. We show the relative improvements and losses in Table 7 vs. a zero-shot baseline on PaLM 2-L-Unicorn. We see gains for 4 out of 5 of the tasks. This provides initial indication that it maybe possible to train separate reward model classifiers to assist in backtracking and that these reward models do not have to be large. Further, a reward model can work on mistakes that are out-of-distribution. However, we believe more data may be necessary to improve results across the board on all tasks. We leave the collection of this larger dataset and a more rigorous investigation of the trade-offs of model size vs. performance of the reward model to future work.
We also leave for future work the effect of backtracking iteratively with a reward model: for example, the generator model may make another mistake after backtracking for the first time, which can then be identified and corrected again.
5 Related work
Datasets
To our knowledge, the only publicly available dataset containing mistake annotations in LLM outputs is PRM800K (Lightman et al., 2023), which is a dataset of solutions to Olympiad-level math questions. Our dataset BIG-Bench Mistake covers a wider range of tasks to explore the reasoning capabilities of LLMs more thoroughly. Additionally, the generator LLM used in PRM800K has been fine-tuned on 1.5B math tokens as well as a dataset of step-by-step math solutions. For this paper, we wanted to explore few-shot in-context learning methods, which is typically used in real-world applications with API-based LLMs.
Self-correction
Pan et al. (2023) present a plethora of self-correction methods in recent literature. While their list includes training-time correction strategies such as RLHF (Ouyang et al., 2022) and self-improve (Huang et al., 2022), our backtracking method falls into the category of post-hoc correction, where the correction process is applied to outputs that have already been generated.
Our paper focuses on correction of logical and reasoning errors, rather than stylistic or qualitative improvements. Previous post-hoc correction methods that are applied to reasoning errors include Reflexion (Shinn et al., 2023) and RCI (Kim et al., 2023), both of which cause performance deterioration when the oracle label is not used (Huang et al., 2023). Other methods such as Self-Refine (Madaan et al., 2023) and iterative refinement (Chen et al., 2023) focus on qualitative or stylistic improvements rather than correcting logical errors.
6 Conclusion
In this paper, we describe and release our dataset BIG-Bench Mistake for mistake finding, and propose a backtracking method to correct logical errors in CoT style traces. We show that LLMs generally struggle with finding logical errors without external feedback, but argue that this feedback can come from a reward model instead. Finally, we demonstrate the effectiveness of backtracking, both with gold standard labels as well as with simulated reward models at lower levels of accuracy.
Limitations
One main limitation of our dataset is that it features tasks that are artificial and unrealistic for real-world applications. We made this choice to minimise ambiguity and subjectivity during the mistake finding process, but further work needs to be done to determine the effectiveness of backtracking in a more realistic setting.
Another limitation is that our paper does not evaluate backtracking on the original datasets on BIG-Bench, only showing results on the limited set that we sampled in a skewed manner, in order to maximise the value of the human annotators’ time. We leave the full evaluation to future work.
Appendix A Implementational details
A.1 3-shot CoT prompting to generate traces for BIG-Bench Mistake
We use PaLM 2 L (Unicorn) to generate the traces used in BIG-Bench Mistake. All traces are generated at temperature .
Our prompts and examples can be found at https://github.com/WHGTyen/BIG-Bench-Mistake. Our prompts are based on chain-of-thought prompts in the BIG-Bench Hard dataset (Suzgun et al., 2022), with four main changes:
1.
Example CoT traces in the prompt is broken up into smaller steps (typically one sentence per step). This is done so that mistake location information is more precise.
2.
Following Yao et al. (2022), each step in the prompt is signposted with “Thought 1”, “Thought 2:”, etc. This allows us to refer to the number of the step when prompting for mistake location.
3.
For the logical deduction task, we find that the notation used in the original prompt with question marks is often inconsistent. It becomes difficult for annotators to determine whether a question mark is a mistake or not, because the correctness of the question mark is dependent on its interpretation. To minimise such ambiguity, the question mark notation is rewritten into text.
4.
For the multistep arithmetic task, one of the prompt examples is altered to increase the length of the equation. This is because the BIG-Bench Hard dataset (where the prompts are taken from) only used equations of a specific length, but our dataset contains equations of averaged a variety of lengths, in accordance with the original BIG-Bench dataset (Srivastava et al., 2022).
Following Yao et al. (2022), we use the newline as the stop token, which generates one step with every generation call. We algorithmically append “Thought N:” before each step. This allows us to split up steps in a clear and systematic way. We stop generating once an answer is reached, which is detected using the following regex:
(?<=[Tt]he answer is).*$
A.2 3-shot prompting to identify mistakes in BIG-Bench Mistake
As described in section 3, we explore three different methods of prompting for mistake location: direct trace-level prompting, direct step-level prompting, and CoT step-level prompting. We use 3-shot prompting for all methods, and our prompts and examples can be found at https://github.com/WHGTyen/BIG-Bench-Mistake.
Our prompts follow OpenAI’s chat completion format. All results were obtained with temperature and no stop tokens.
Appendix B Annotation
We release our annotation guidelines at https://github.com/WHGTyen/BIG-Bench-Mistake.
During annotation of the multistep arithmetic task, we found that the first CoT step given in the original BIG-Bench Hard prompt examples (Suzgun et al., 2022) was incorrect. Since all generated traces contained the same first step, we removed that step before showing traces to the annotators.
Figure 3 contains an example screenshot of the user interface. For every trace, we provide the input question as well as the target answer, with a note to be aware of errors that may occur in correct traces.
Annotators can click on words to highlight the same word across the trace and the question text, which we found was particularly helpful for some tasks such as word sorting and tracking shuffled objects. Buttons on the right automatically become inactive if a previous step has been labelled as negative.
| Model |
Direct (trace) |
Direct (step) |
CoT (step) |
| Word sorting | |||
| GPT-4-Turbo |
67.74 |
38.24 |
| GPT-4 |
88.24 |
82.35 |
58.82 |
| GPT-3.5-Turbo |
100.00 |
97.06 |
20.59 |
| Tracking shuffled objects | |||
| GPT-4-Turbo |
90.00 |
77.50 |
| GPT-4 |
82.50 |
82.50 |
80.00 |
| GPT-3.5-Turbo |
67.50 |
0.00 |
0.00 |
| Logical deduction | |||
| GPT-4-Turbo |
100.00 |
83.33 |
| GPT-4 |
100.00 |
100.00 |
0.00 |
| GPT-3.5-Turbo |
100.00 |
50.00 |
100.00 |
| Multistep arithmetic | |||
| GPT-4-Turbo |
57.69 |
40.32 |
| GPT-4 |
53.23 |
46.77 |
27.42 |
| GPT-3.5-Turbo |
96.77 |
79.03 |
58.06 |
| Dyck languages | |||
| GPT-4-Turbo |
96.42 |
30.00 |
| GPT-4 |
98.41 |
78.57 |
13.79 |
| GPT-3.5-Turbo |
95.74 |
4.76 |
0.00 |
| Model |
Direct (trace) |
Direct (step) |
CoT (step) |
| Word sorting | |||
| GPT-4-Turbo |
32.71 |
32.33 |
| GPT-4 |
28.20 |
39.47 |
30.83 |
| GPT-3.5-Turbo |
0.00 |
4.51 |
15.04 |
| Tracking shuffled objects | |||
| GPT-4-Turbo |
31.54 |
59.23 |
| GPT-4 |
59.14 |
62.69 |
92.31 |
| GPT-3.5-Turbo |
1.17 |
1.92 |
21.92 |
| Logical deduction | |||
| GPT-4-Turbo |
20.81 |
74.83 |
| GPT-4 |
39.46 |
67.01 |
10.54 |
| GPT-3.5-Turbo |
0.00 |
24.83 |
7.82 |
| Multistep arithmetic | |||
| GPT-4-Turbo |
34.27 |
44.12 |
| GPT-4 |
41.60 |
41.60 |
44.54 |
| GPT-3.5-Turbo |
0.00 |
12.18 |
16.81 |
| Dyck languages | |||
| GPT-4-Turbo |
6.99 |
28.46 |
| GPT-4 |
7.37 |
40.81 |
43.91 |
| GPT-3.5-Turbo |
1.28 |
6.05 |
2.08 |

### v3

@@
-LLMs cannot find reasoning errors, but can correct them!
+LLMs cannot find reasoning errors, but can correct them given the error location
 Abstract
 While self-correction has shown promise in improving LLM outputs in terms of style and quality (e.g.
-Chen et al., 2023; Madaan et al., 2023), recent attempts to self-correct logical or reasoning errors often cause correct answers to become incorrect, resulting in worse performances overall (Huang et al., 2023).
-In this paper, we break down the self-correction process into two core components: mistake finding and output correction.
-For mistake finding, we release BIG-Bench Mistake, a dataset of logical mistakes in Chain-of-Thought reasoning traces.
-We provide benchmark numbers for several state-of-the-art LLMs, and demonstrate that LLMs generally struggle with finding logical mistakes.
-For output correction, we propose a backtracking method which provides large improvements when given information on mistake location.
-We construe backtracking as a lightweight alternative to reinforcement learning methods, and show that it remains effective with a reward model at 60-70% accuracy.
+Chen et al., 2023b; Madaan et al., 2023), recent attempts to self-correct logical or reasoning errors often cause correct answers to become incorrect, resulting in worse performances overall (Huang et al., 2023).
+In this paper, we show that poor self-correction performance stems from LLMs’ inability to find logical mistakes, rather than their ability to correct a known mistake.
+Firstly, we benchmark several state-of-the-art LLMs on their mistake-finding ability and demonstrate that they generally struggle with the task, even in highly objective, unambiguous cases.
+Secondly, we test the correction abilities of LLMs – separately from mistake finding – using a backtracking setup that feeds ground truth mistake location information to the model.
+We show that this boosts downstream task performance across our 5 reasoning tasks, indicating that LLMs’ correction abilities are robust.
+Finally, we show that it is possible to obtain mistake location information without ground truth labels or in-domain training data.
+We train a small classifier with out-of-domain data, which exhibits stronger mistake-finding performance than prompting a large model.
+We release our dataset of LLM-generated logical mistakes, BIG-Bench Mistake, to enable further research into locating LLM reasoning mistakes.
 1 Introduction
@@
 Recent literature on few- or zero-shot prompting has focused on the concept of self-correction, i.e.
-having an LLM correct its own outputs (Shinn et al., 2023; Miao et al., 2023; Madaan et al., 2023; Chen et al., 2023; Saunders et al., 2022).
+having an LLM correct its own outputs (Shinn et al., 2023; Miao et al., 2024; Madaan et al., 2023; Chen et al., 2023b; Saunders et al., 2022).
 (See Pan et al.
@@
 However, Huang et al.
-(2023) note that while self-correction may prove effective for improving model outputs in terms of style and quality, there is limited evidence that LLMs can identify and fix their own reasoning and logical errors without external feedback.
-For example, Reflexion (Shinn et al., 2023) and RCI (Kim et al., 2023) both use ground truth correctness as a signal to halt the self-correction loop.
+(2023) note that while self-correction may prove effective for improving model outputs in terms of style and quality, when it comes to reasoning tasks, LLMs struggle to identify and fix errors without external feedback: for example, Reflexion (Shinn et al., 2023) and RCI (Kim et al., 2023) both use ground truth correctness as a signal to halt the self-correction loop.
 Initially observed by Madaan et al.
-(2023) on a math reasoning dataset, Huang et al.
+(2023) on a math dataset, Huang et al.
 (2023) further demonstrate this shortcoming of self-correction in 2 additional datasets.
-While previous work typically present self-correction as a single process, we divide it into mistake finding and output correction.
+While previous work typically present self-correction as a single process, we divide it into mistake finding and output correction to better understand each component individually.
 Mistake finding is a fundamental reasoning skill that has been studied and utilised extensively in philosophy, psychology, and mathematics, spawning concepts such as critical thinking, and logical and mathematical fallacies.
@@
 Output correction involves partially or completely changing previously generated outputs.
-In the context of self-correction, this is typically done with outputs generated by the same model (see Pan et al.
-(2023) for an overview of different strategies).
-Despite LLMs’ inability to find mistakes, our results show that they can correct outputs using our backtracking method, if given information about the mistakes, for example via a small, supervised reward model.
+With self-correction, this is typically done with outputs generated by the same model (see Pan et al.
+(2023)).
+Despite LLMs’ inability to find mistakes, our results show that they can correct outputs, if given information about the mistake location.
+While LLMs struggle with mistake-finding in few-shot conditions, we can obtain more reliable mistake location information using a small, trained classifier.
 Our contributions for this paper are as follows:
@@
 With Chain-of-Thought prompting, any task can be turned into a mistake-finding task.
-We collect and release111https://github.com/WHGTyen/BIG-Bench-Mistake.
-to the research community BIG-Bench Mistake, a dataset of CoT-style traces222In this paper, we refer to a set of CoT reasoning steps as a trace.
-generated using PaLM 2, and annotated according to where the first logical mistake is.
+We collect and release111Publicly available at https://github.com/WHGTyen/BIG-Bench-Mistake.
+to the research community BIG-Bench Mistake, a dataset of CoT-style traces222We refer to a set of CoT reasoning steps as a trace.
+generated using PaLM 2 (Anil et al., 2023), and annotated according to where the first logical mistake is.
 To our knowledge, BIG-Bench Mistake is the first dataset of its kind that goes beyond problems in mathematics.
 2.
-We produce benchmark results for our dataset to test the reasoning capabilities of state-of-the-art LLMs.
-We demonstrate that current state-of-the-art LLMs struggle with mistake finding, even for objective, unambiguous cases.
-We hypothesise that this is a main contributing factor to LLMs’ inability to self-correct reasoning errors, and we call on the research community to pursue further improvements on the mistake finding task.
+We produce benchmark results for our dataset to test the reasoning capabilities of five state-of-the-art LLMs.
+We demonstrate that these LLMs struggle with mistake finding, even for objective, unambiguous cases.
+We hypothesise that this is a main contributing factor to LLMs’ inability to self-correct reasoning errors.
 3.
-We propose backtracking as an output correction technique that makes use of mistake location information to improve performance on the original task.
-We demonstrate that this method corrects outputs that are originally incorrect, with minimal effect on outputs that are originally correct.
+We test LLMs’ ability to correct reasoning errors separately from mistake-finding, by feeding to the model the ground truth (or oracle) mistake location information through a backtracking method.
+We demonstrate that LLMs’ correction abilities are robust, effectively correcting outputs that are originally incorrect, with minimal effect on outputs that are originally correct.
 4.
-We construe backtracking as a form of “verbal reinforcement learning” (Shinn et al., 2023), allowing iterative improvement on CoT outputs without requiring any weight updates.
-We propose that backtracking can be used with a trained classifier as a reward model, and demonstrate the effectiveness of backtracking at various reward model accuracies.
+We demonstrate that LLMs’ low accuracy at few-shot mistake-finding can be improved upon without using in-domain training data.
+As a proof-of-concept, we train a small mistake-finding classifier using out-of-domain data, which performs better than prompting a large model.
+We leave the development of more sophisticated methods to future work.
 2 BIG-Bench Mistake
-BIG-Bench Mistake consists of 2186 sets of CoT-style traces.
-Each trace was generated by PaLM 2-L-Unicorn, and annotated with the location of the first logical error.
-An example trace is shown in Table 1, where the mistake location333As some traces may not contain mistakes, we use the term mistake location as a multi-class label that can refer to either the integer where the th step contains the first mistake, or that there are no mistakes.
+BIG-Bench Mistake contains 2186 sets of CoT-style traces.
+Each trace is generated with PaLM 2 Unicorn (Anil et al., 2023), and annotated with the location of the first logical error.
+Table 1 shows an example trace where the mistake location333As some traces may not contain mistakes, we use the term mistake location as a multi-class label that can refer to either the integer where the th step contains the first mistake, or that there are no mistakes.
 is the 4th step.
 Sort the following words alphabetically: List: hypochlorite ponderosa phone credulity | |
-| Thought 1: |
-I should start by looking at the first letter of the words in the list. The first letter: "hypochlorite": "h" (8). "ponderosa": "p" (16). "phone": "p" (16). "credulity": "c" (3). |
-| Thought 2: |
-We now have: (3) "credulity" < (8) "hypochlorite" < (16) ["ponderosa" ? "phone"]. |
-| Thought 3: |
-Now let’s sort this subpart ["ponderosa" ? "phone"] by looking at their second letters. The second letter: "ponderosa": "o" (15). "phone": "h" (8). |
-| Thought 4: |
-We now have: (8) "phone" < (15) "ponderosa" for the subpart. Hence, we have "credulity" < "phone" < "ponderosa". (MISTAKE) |
-| Thought 5: |
-I have now sorted all the words. The answer is credulity hypochlorite phone ponderosa |
-Our traces span across a set of 5 tasks444These 5 tasks were selected because 1) Anil et al.
-(2023) demonstrate that PaLM 2 performs poorly on these tasks, so it is likely to generate mistakes in CoT traces; 2) any mistakes that may occur are likely to be unambiguous, therefore minimising subjectivity during annotation; and 3) identifying mistakes for these tasks does not require expertise knowledge of a specific domain.
-from the BIG-bench dataset (Srivastava et al., 2023): word sorting, tracking shuffled objects, logical deduction, multi-step arithmetic, and Dyck languages.
+Thought 1: | I should start by looking at the first letter of the words in the list. The first letter: "hypochlorite": "h" (8). "ponderosa": "p" (16). "phone": "p" (16). "credulity": "c" (3). |
+Thought 2: | We now have: (3) "credulity" < (8) "hypochlorite" < (16) ["ponderosa" ? "phone"]. |
+Thought 3: | Now let’s sort this subpart ["ponderosa" ? "phone"] by looking at their second letters. The second letter: "ponderosa": "o" (15). "phone": "h" (8). |
+Thought 4:
+We now have: (8) "phone" < (15) "ponderosa" for the subpart. Hence, we have "credulity" < "phone" < "ponderosa". |
+Thought 5: | I have now sorted all the words. The answer is credulity hypochlorite phone ponderosa |
+Our traces span across a set of 5 tasks from the BIG-bench dataset (Srivastava et al., 2023): word sorting, tracking shuffled objects, logical deduction, multi-step arithmetic, and Dyck languages444These 5 tasks are selected because 1) Anil et al.
+(2023) demonstrate that PaLM 2 performs poorly on these tasks, so it is likely to generate mistakes in CoT traces; 2) mistakes in these tasks are likely to be unambiguous, therefore minimising subjectivity during annotation; and 3) identifying mistakes for these tasks does not require expertise knowledge..
 CoT prompting is used to prompt PaLM 2 to answer questions from each task.
-As we wanted to separate our CoT traces into distinct steps, we follow the method used by Yao et al.
+As we wanted to separate our CoT traces into distinct steps, we follow Yao et al.
 (2022) and generate each step separately, using the newline as a stop token.
-In this dataset, all traces are generated with temperature . The correctness of answers are determined by exact match. Prompts can be found at https://github.com/WHGTyen/BIG-Bench-Mistake.
+All traces are generated with temperature . The correctness of answers are determined by exact match. Prompts can be found at https://github.com/WHGTyen/BIG-Bench-Mistake along with the dataset.
+Task | # of correctans traces | # of incorrectans traces | # of incorrectmis traces | Total |
+Word sorting | 45 | 255 | 266 | 300 |
+Tracking shuffled objects | 45 | 255 | 260 | 300 |
+Logical deduction | 45 | 255 | 294 | 300 |
+Multistep arithmetic | 45 | 255 | 238 | 300 |
+Dyck languages | 482 | 504 | 650 | 986 |
+Dyck languages (sampled) | 88 | 504 | 545 | 592 |
 2.1 Annotation
@@
 Note that traces can contain a logical mistake yet arrive at the correct answer.
-To disambiguate the two types of correctness, we will use the terms correct and incorrect to refer to whether the final answer of the trace is correct.
-Accuracy would therefore refer to the overall accuracy for the task, based on how many final answers are correct.
-To refer to whether the trace contains a logical mistake (rather than the correctness of the final answer), we will use correct and incorrect.
+To disambiguate the two types of correctness, we will use the terms correctans and incorrectans to refer to whether the final answer of the trace is correct.
+Accuracyans would therefore refer to the overall accuracy for the task, based on how many final answers are correct.
+To refer to whether the trace contains a logical mistake (rather than the correctness of the final answer), we will use correctmis and incorrectmis.
 2.1.1 Human annotation
 For 4 of the 5 tasks, we recruit human annotators to go through each trace and identify any errors.
-Annotators have no domain expertise but are given guidelines555See https://github.com/WHGTyen/BIG-Bench-Mistake for further details.
+Annotators have no domain expertise but are given guidelines555https://github.com/WHGTyen/BIG-Bench-Mistake contains further details.
 to complete the task.
-Before annotation, we sample a set of 300 traces for each task, where 255 (85%) are incorrect, and 45 (15%) are correct.
+Before annotation, we sample a set of 300 traces for each task, where 255 (85%) are incorrectans, and 45 (15%) are correctans.
 Since human annotation is a limited and expensive resource, we chose this distribution to maximise the number of steps containing mistakes and to prevent over-saturation of correct steps.
-We also include some correct traces because some may contain logical errors despite the correct answer, and to ensure that the dataset included examples of correct steps that are near the end of the trace.
-This also prevents annotators from feeling forced to find a mistake in all traces.To account for this skewed distribution, results in section 4 are split according to whether the original trace is correct or incorrect.
+We also include some correctans traces because some may contain logical errors despite the correct answer, and to ensure that the dataset included examples of correct steps that are near the end of the trace.
+To account for this skewed distribution, results in section 4 are split according to whether the original trace is correctans or not.
 Following Lightman et al.
 (2023), annotators are instructed to go through each step in the trace and indicate whether the step is correct or not (binary choice).
-Annotators only submit their answers until all steps have been annotated, or there is one incorrect step.
+Annotators only submit their answers when all steps are annotated, or there is one incorrect step.
 If an incorrect step is identified, the remaining steps are skipped.
-This is done to avoid ambiguities where a logically correct deduction is dependent on a previous mistake.
-We make our annotation guidelines available at https://github.com/WHGTyen/BIG-Bench-Mistake, and we include a screenshot of the user interface in Figure 3.
-Each trace has been annotated by at least 3 annotators.
+This is to avoid ambiguities where a logically correct deduction is dependent on a previous mistake.
+Our annotation guidelines can be found at https://github.com/WHGTyen/BIG-Bench-Mistake/tree/main/annotation_guidelines, and we include a screenshot of the user interface in Appendix D.
+Each trace is annotated by at least 3 annotators.
 If there are any disagreements, we take the majority label.
@@
 2.1.2 Automatic annotation
-For Dyck languages, we opt for mostly automatic annotation instead of human annotation as the traces show limited variation in phrasing and solution paths.
-For each trace, we generate a set of standard steps based on the format used in the prompt examples.
-Using pattern matching, we can identify whether each model-generated step also conforms to the same format.
+For Dyck languages, we use mostly automatic instead of human annotation, as the traces show limited variation in phrasing and solution paths.
+For each trace, we algorithmically generate a set of steps based on the format used in the prompt examples.
+Using pattern matching, we identify whether each model-generated step conforms to the same format.
 If so, we compare the two and assume that the trace is incorrect if the symbols do not match.
-Additionally, we also account for edge cases such as where the final two steps are merged into one, or variations in presentation where some symbols are placed in quotes and some are not.
+Additionally, we account for edge cases such as where the final two steps are merged into one, or variations in presentation where symbols may or may not be placed in quotes.
 We release the code at https://github.com/WHGTyen/BIG-Bench-Mistake along with our dataset.
-| Task |
-Num. of correct traces |
-Num. of incorrect |
-Num. of incorrect traces |
-Total |
-| Word sorting |
-45 |
-255 |
-266 |
-300 |
-| Tracking shuffled objects |
-45 |
-255 |
-260 |
-300 |
-| Logical deduction |
-45 |
-255 |
-294 |
-300 |
-| Multistep arithmetic |
-45 |
-255 |
-238 |
-300 |
-| Dyck languages |
-482 |
-504 |
-650 |
-986 |
-| Dyck languages (sampled) |
-88 |
-504 |
-545 |
-592 |
-3 Benchmark results
-Table 4 shows the accuracy of GPT-4-Turbo, GPT-4, and GPT-3.5-Turbo on our mistake-finding dataset.
-For each question, the possible answers are either that there are no mistakes, or, if there is a mistake, the number N indicating the step in which the first mistake occurs.
+3 Can LLMs find reasoning mistakes in CoT traces?
+Table 4 shows the accuracy of GPT-4-Turbo, GPT-4, GPT-3.5-Turbo, Gemini Pro, and PaLM 2 Unicorn on our mistake-finding dataset.
+For each question, the possible answers are either: that there are no mistakes, or; if there is a mistake, the number N indicating the step in which the first mistake occurs.
 A model’s output is only considered correct if the location matches exactly, or the output correctly indicates that there are no mistakes.
-All models are given the same 3-shot prompts666Prompts can be found at https://github.com/WHGTyen/BIG-Bench-Mistake.. We use three different prompting methods:
+All models are given the same 3-shot promptsfootnote 5. We use three different prompting methods:
 Direct trace-level prompting involves using the whole trace as input to the model and directly prompting for the mistake location.
@@
 This method is the most resource intensive of all three methods as it involves generating a whole CoT sequence for every step.
-Due to cost and usage limits, we are unable to provide results from GPT-4-Turbo here.
 As with direct step-level prompting, the final answer is inferred from where the first "No" output occurs (subsequent steps are ignored).
 3.1 Discussion
-All three models appear to struggle with our mistake finding dataset. GPT-4 attains the best results but only reaches an overall accuracy of 52.87 with direct step-level prompting.
+All five models appear to struggle with our mistake finding dataset.
+GPT-4 attains the best results but only reaches an overall accuracy of 52.87 with direct step-level prompting.
+While exact parameter counts are undisclosed, GPT-4 is likely one of the largest models, along with PaLM 2 Unicorn666Note that the traces in our dataset are generated using PaLM 2 Unicorn and are sampled according to whether the final answer was correct or not.
+Therefore, we expect that using PaLM 2 itself to do mistake finding will produce different and likely biased results.
+Further work is needed to elucidate the difference between cross-model evaluation and self-evaluation., while Gemini Pro and GPT-3.5-Turbo are among the smaller models.
 Our findings are in line with and builds upon results from Huang et al.
@@
 If LLMs are unable to identify mistakes, it should be no surprise that they are unable to self-correct either.
-Note that the mistakes in our dataset are generated using PaLM 2 L (Unicorn), and traces were sampled according to whether the final answer was correct or not.
-Therefore, we expect that using PaLM 2 itself to do mistake finding will produce different and likely biased results.
-Further work is needed to elucidate the difference between cross-model evaluation and self-evaluation.
-| Model |
-Direct (trace) |
-Direct (step) |
-CoT (step) |
+Model | Direct (trace) | Direct (step) | CoT (step) |
 Word sorting (11.7) | |||
-| GPT-4-Turbo |
-36.33 |
-33.00 |
-| GPT-4 |
-35.00 |
-44.33 |
-34.00 |
-| GPT-3.5-Turbo |
-11.33 |
-15.00 |
-15.67 |
+GPT-4-Turbo | 36.33 | 33.00 | – |
+GPT-4 | 35.00 | 44.33 | 34.00 |
+GPT-3.5-Turbo | 11.33 | 15.00 | 15.67 |
+Gemini Pro | 10.67 | – | – |
+PaLM 2 Unicorn | 11.67 | 16.33 | 14.00 |
 Tracking shuffled objects (5.4) | |||
-| GPT-4-Turbo |
-39.33 |
-61.67 |
-| GPT-4 |
-62.29 |
-65.33 |
-90.67 |
-| GPT-3.5-Turbo |
-10.10 |
-1.67 |
-19.00 |
+GPT-4-Turbo | 39.33 | 61.67 | – |
+GPT-4 | 62.29 | 65.33 | 90.67 |
+GPT-3.5-Turbo | 10.10 | 1.67 | 19.00 |
+Gemini Pro | 37.67 | – | – |
+PaLM 2 Unicorn | 18.00 | 28.00 | 55.67 |
 Logical deduction (8.3) | |||
-| GPT-4-Turbo |
-21.33 |
-75.00 |
-| GPT-4 |
-40.67 |
-67.67 |
-10.33 |
-| GPT-3.5-Turbo |
-2.00 |
-25.33 |
-9.67 |
+GPT-4-Turbo | 21.33 | 75.00 | – |
+GPT-4 | 40.67 | 67.67 | 10.33 |
+GPT-3.5-Turbo | 2.00 | 25.33 | 9.67 |
+Gemini Pro | 8.67 | – | – |
+PaLM 2 Unicorn | 6.67 | 38.00 | 12.00 |
 Multistep arithmetic (5.0) | |||
-| GPT-4-Turbo |
-38.33 |
-43.33 |
-| GPT-4 |
-44.00 |
-42.67 |
-41.00 |
-| GPT-3.5-Turbo |
-20.00 |
-26.00 |
-25.33 |
+GPT-4-Turbo | 38.33 | 43.33 | – |
+GPT-4 | 44.00 | 42.67 | 41.00 |
+GPT-3.5-Turbo | 20.00 | 26.00 | 25.33 |
+Gemini Pro | 21.67 | – | – |
+PaLM 2 Unicorn | 22.00 | 21.67 | 23.67 |
 Dyck languages† (24.5) | |||
-| GPT-4-Turbo |
-15.33* |
-28.67* |
-| GPT-4 |
-17.06 |
-44.33* |
-41.00* |
-| GPT-3.5-Turbo |
-8.78 |
-5.91 |
-1.86 |
+GPT-4-Turbo | 15.33* | 28.67* | – |
+GPT-4 | 17.06 | 44.33* | 41.00* |
+GPT-3.5-Turbo | 8.78 | 5.91 | 1.86 |
+Gemini Pro | 2.00 | – | – |
+PaLM 2 Unicorn | 10.98 | 14.36 | 17.91 |
 Overall | |||
-| GPT-4-Turbo |
-30.13 |
-48.33 |
-| GPT-4 |
-39.80 |
-52.87 |
-43.40 |
-| GPT-3.5-Turbo |
-10.44 |
-14.78 |
-14.31 |
-† indicates that traces were sampled to contain 15% correct and 85% incorrect traces (see Table 3).
-* indicates that traces were sampled to contain 45 correct and 255 incorrect traces to reduce costs.
+GPT-4-Turbo | 30.13 | 48.33 | – |
+GPT-4 | 39.80 | 52.87 | 43.40 |
+GPT-3.5-Turbo | 10.44 | 14.78 | 14.31 |
+Gemini Pro | 16.14 | – | – |
+PaLM 2 Unicorn | 17.09 | 23.67 | 24.65 |
+† indicates that traces were sampled to contain 15% correctans and 85% incorrectans traces (see Table 2).
+* indicates that traces were sampled to contain 45 correctans and 255 incorrectans traces to reduce costs.
 3.2 Comparison of prompting methods
-As we compare results across the three methods, we find that the accuracy on traces with no mistakes goes down777Note that the traces in BIG-Bench Mistake are sampled to contain more incorrect traces than correct traces (and therefore more incorrect traces than correct traces), so the overall mistake location accuracy appears higher for per-step prompting in Table 4, despite the poor accuracy for correct traces.
-For a full set of split by correctness, see Appendix B.
+As we compare results across the three methods, we find that the accuracy on traces with no mistakes goes down777Note that the traces in BIG-Bench Mistake are sampled to contain more incorrectans traces than correctans traces (and therefore more incorrectmis traces than correctmis traces), so the overall mistake location accuracy appears higher for per-step prompting in Table 4, despite the poor accuracy for correctmis traces.
+For a full set of scores split by correctnessmis, see Appendix E.
 considerably from direct, trace-level prompting to CoT, step-level prompting.
@@
 3.3 Few-shot prompting for mistake location as a proxy for correctness
-In this section, we investigate whether our prompting methods can reliably determine the correctness of a trace rather than the mistake location.
-Our motivation was that even humans use mistake finding as a strategy for determining whether an answer is correct or not, such as when going through mathematical proofs, or working through argumentation.
-Additionally, one might think that directly predicting the correctness of a trace may be easier than having to pinpoint the precise location of an error.
-We calculate averaged F1 scores based on whether the model predicts that there is a mistake in the trace.
-If there is a mistake, we assume the model prediction is that the trace is incorrect.
-Otherwise, we assume the model prediction is that the trace is correct.
-| Model |
-Direct (trace) |
-Direct (step) |
-CoT (step) |
+In this section, we investigate whether our prompting methods can reliably determine the correctnessans of a trace rather than the mistake location.
+Our motivation was that even humans use mistake finding as a strategy for determining whether an answer is correct or not, like when going through mathematical proofs or argumentation.
+Additionally, it may be the case that directly predicting the correctnessans of a trace is easier than pinpointing the precise location of an error.
+Model | Direct (trace) | Direct (step) | CoT (step) |
 Word sorting | |||
-| GPT-4-Turbo |
-87.73 |
-86.68 |
-| GPT-4 |
-81.50 |
-85.12 |
-81.19 |
-| GPT-3.5-Turbo |
-6.58 |
-35.07 |
-77.79 |
+GPT-4-Turbo | 87.73 | 86.68 | – |
+GPT-4 | 81.50 | 85.12 | 81.19 |
+GPT-3.5-Turbo | 6.58 | 35.07 | 77.79 |
+Gemini Pro | 69.93 | – | – |
+PaLM 2 Unicorn | 21.08 | 56.66 | 62.92 |
 Tracking shuffled objects | |||
-| GPT-4-Turbo |
-52.23 |
-74.31 |
-| GPT-4 |
-76.38 |
-75.69 |
-95.03 |
-| GPT-3.5-Turbo |
-32.04 |
-77.61 |
-78.11 |
+GPT-4-Turbo | 52.23 | 74.31 | – |
+GPT-4 | 76.38 | 75.69 | 95.03 |
+GPT-3.5-Turbo | 32.04 | 77.61 | 78.11 |
+Gemini Pro | 79.66 | – | – |
+PaLM 2 Unicorn | 22.18 | 48.77 | 78.29 |
 Logical deduction | |||
-| GPT-4-Turbo |
-86.46 |
-81.79 |
-| GPT-4 |
-84.54 |
-83.38 |
-23.96 |
-| GPT-3.5-Turbo |
-10.34 |
-67.62 |
-61.31 |
+GPT-4-Turbo | 86.46 | 81.79 | – |
+GPT-4 | 84.54 | 83.38 | 23.96 |
+GPT-3.5-Turbo | 10.34 | 67.62 | 61.31 |
+Gemini Pro | 48.57 | – | – |
+PaLM 2 Unicorn | 31.67 | 37.93 | 21.21 |
 Multistep arithmetic | |||
-| GPT-4-Turbo |
-71.17 |
-86.24 |
-| GPT-4 |
-72.97 |
-78.67 |
-79.67 |
-| GPT-3.5-Turbo |
-3.76 |
-53.18 |
-64.08 |
+GPT-4-Turbo | 71.17 | 86.24 | – |
+GPT-4 | 72.97 | 78.67 | 79.67 |
+GPT-3.5-Turbo | 3.76 | 53.18 | 64.08 |
+Gemini Pro | 32.21 | – | – |
+PaLM 2 Unicorn | 33.69 | 13.42 | 70.94 |
 Dyck languages | |||
-| GPT-4-Turbo |
-51.96 |
-85.87 |
-| GPT-4 |
-62.33 |
-85.73 |
-79.60 |
-| GPT-3.5-Turbo |
-46.57 |
-79.31 |
-77.79 |
-In Table 5, we average the F1s using correct and incorrect as the positive label, weighted according to the number of times each label occurs.
-Note that the baseline of predicting all traces as incorrect achieves a weighted F1 average of 78.
-The weighted F1 scores show that prompting for mistakes is a poor strategy for determining the correctness of the final answer.
+GPT-4-Turbo | 51.96 | 85.87 | – |
+GPT-4 | 62.33 | 85.73 | 79.60 |
+GPT-3.5-Turbo | 46.57 | 79.31 | 77.79 |
+Gemini Pro | 61.24 | – | – |
+PaLM 2 Unicorn | 31.17 | 31.63 | 25.20 |
+We calculate averaged F1 scores based on whether the model predicts there is a mistake in the trace.
+If there is a mistake, we assume the model prediction is that the trace is incorrectans.
+Otherwise, we assume the model prediction is that the trace is correctans.
+In Table 5, we average the F1s calculated with correctans and incorrectans as positive labels, weighted according to the number of times each label occurs.
+Note that the naive baseline of predicting all traces as incorrect achieves a weighted F1 average of 78.
+The weighted F1 scores show that prompting for mistakes is likely a poor strategy for determining the correctness of the final answer.
 This is in line with our previous finding that LLMs struggle to identify mistake locations, and also builds upon results from Huang et al.
-(2023), who demonstrate that improvements from Reflexion (Shinn et al., 2023) and RCI (Kim et al., 2023) are only from using oracle correctness information.
-4 Backtracking
-Madaan et al.
-(2023) and Huang et al.
-(2023) both demonstrate that self-correction is only effective with external feedback – for example, both Shinn et al.
-(2023) and Kim et al.
-(2023) rely on oracle labels for improvements – but there is often no external feedback available in many real-world applications.
-As an alternative, we explore the possibility of replacing external feedback with a lightweight classifier trained on a small amount of data.
-Analogous to reward models in conventional reinforcement learning, this classifier detects any logical errors in a CoT trace, which is then fed back to the generator model to improve on the output.
-This can be done over multiple iterations to maximise improvements.
-We propose a simple backtracking method to improve model outputs based on the location of logical errors:
-1.
+(2023), who demonstrate that improvements from Reflexion (Shinn et al., 2023) and RCI (Kim et al., 2023) are only from using oracle correctnessans information.
+4 Can LLMs correct reasoning mistakes in CoT traces?
+In this section, we examine LLMs’ ability to correct mistakes, independently of their ability to find them.
+To do so, we feed oracle mistake location information from BIG-Bench Mistake into the model and prompt it for a corrected version of the original CoT trace.
+As a simple baseline, we use the following backtracking method (visualized in Figure 2):
 First, the model generates an initial CoT trace. In our experiments, we use temperature .
-2.
-We then determine the mistake location in this trace using a reward model.
-3.
+We then determine the mistake location in this trace, either from oracle labels (in this section) or with a classifier (in section 5).
 If there are no mistakes, we move onto the next trace.
 If there is a mistake (e.g.
-at Thought 4 in the example trace in Table 1),we prompt the model again for the same step but at temperature , generating 8 outputs.
+at Thought 4 in the example trace in Table 1), we prompt the model again for the same step but at temperature .
 We use same prompt and the partial trace containing all steps up to but not including the mistake step (e.g.
 up to Thought 3, prompting for Thought 4).
-4.
-From the 8 outputs, we filter out any options that match what was previously identified as a mistake. From the remaining outputs, we select one with the highest log-probability.
-5.
+In our experiments, we found that (c) often produced steps that are identical to the original.
+We therefore repeat (c) until a different step is generated (or up to a fixed number, whichever is less).
+For this paper, we use 8 as the maximum number of re-generations; the effects of varying this number is left for future investigation.
+To reduce computational cost, we generate 8 outputs simultaneously but only select one for backtracking.
 Finally, with the new, regenerated step in place of the previous one, we generate the remaining steps of the trace again at temperature .
-Our backtracking method provides several benefits over existing self-correction methods:
-Unlike Shinn et al.
-(2023), Kim et al.
-(2023), etc., our approach does not depend on oracle knowledge of the answer.
-Instead, it relies on information (for example from trained a reward model) about logical errors, which can be determined on a step-by-step basis using a reward model.
-Logical errors can occur in correct traces, or not occur in incorrect traces888Having no logical errors in incorrect traces is much rarer but does exist, for example when the answer is correct but is not captured by exact match, or if the original question is faulty and has multiple possible answers..
-Compared to approaches that require re-generating the entire trace, backtracking reduces computational cost by reusing previous steps that are known to be logically sound.
-Backtracking improves on the quality of the intermediate steps directly, which can be useful in scenarios that require correct steps (e.g.
-generating solutions to math questions), and also generally improves interpretability.
-Backtracking with mistake location information from a reward model can be construed as a lightweight RL method. However, unlike conventional deep reinforcement learning:
-Backtracking with a reward model does not does not require any training of the original generator model.
-Once the reward model is trained, it can be used for backtracking with any LLM as the generator, and can also be updated independently of the generator LM.
-This can be especially helpful when LLMs are frequently updated to new checkpoints.
-Backtracking only requires training of a small reward model.
-Compared to methods that require training of the generator model, backtracking is far more efficient in terms of computing resources and available data.
-The process of backtracking is more interpretable than updating the weights of the generator model directly, as is required for many deep RL methods.
-It clearly pinpoints the location at which an error occurs, which can help the debugging process and allow faster development and iterations of models.
-4.1 Backtracking with gold mistake location
-As an initial experiment, we use labels from BIG-Bench Mistake to test if an LLM is able to correct logical errors using backtracking, independent of its inherent ability to identify these errors or any other reward model.
-For example, if the mistake location is in step 4, we use backtracking to regenerate that step and continue the rest of the chain.
-If the mistake location is that there are no logical mistakes, we do not backtrack and use the original result.
-| With mistake location | With random location |
-Avg. num. of steps |
+This backtracking method is designed to be a very simple baseline, with no specific prompt text or phrasing, and without relying on generating a large number of alternatives.
+For our experimental results below, we specifically use the same model (PaLM 2 Unicorn) to correct the traces it originally generated, to test its ability to self-correct.
+With mistake location | With random location | Avg. num. of steps | |||
 Task | accuracy | accuracy | accuracy | accuracy | |
-| Word sorting | -11.11 | +23.53 | -15.56 | +11.76 |
-11.7 |
-| Tracking shuffled objects | -6.67 | +43.92 | -6.67 | +20.39 |
-5.4 |
-| Logical deduction | -11.43 | +36.86 | -13.33 | +21.57 |
-8.3 |
-| Multistep arithmetic | -0.00 | +18.04 | -8.89 | +10.59 |
-5.0 |
-| Dyck languages | -6.82 | +18.06 | -15.91 | +5.16 |
-24.5 |
-4.1.1 Results
+Word sorting | -11.11 | +23.53 | -15.56 | +11.76 | 11.7 |
+Tracking shuffled objects | -6.67 | +43.92 | -6.67 | +20.39 | 5.4 |
+Logical deduction | -11.43 | +36.86 | -13.33 | +21.57 | 8.3 |
+Multistep arithmetic | -0.00 | +18.04 | -8.89 | +10.59 | 5.0 |
+Dyck languages | -6.82 | +18.06 | -15.91 | +5.16 | 24.5 |
+4.1 Results
 The results are shown in Table 6.
-To show that performance increases are not due to randomly resampling outputs, we compare our results to a random baseline, where a mistake location999As described above, the mistake location can be either the number representing the step, or that there are no mistakes.
+To show that performance increases are not due to randomly resampling outputs, we compare our results to a random baseline, where a mistake location888As described above, the mistake location can be either the number representing the step, or that there are no mistakes.
 If there are no mistakes, we do not use backtracking and simply use the original trace.
 is randomly selected for each trace and we perform backtracking based on the random location.
-Note that Table 6 separates results into numbers for the correct set and the incorrect set, referring to whether the original trace was correct or not.
-This gives a clearer picture than the overall accuracy, which would be skewed by the proportion of traces that were originally correct (15%) and incorrect (85%).
-Scores represent the absolute differences in accuracy. We perform backtracking on both correct and incorrect traces, as long as there is a mistake in one of the steps.
+Note that Table 6 separates results into numbers for the correct set and the incorrect set, referring to whether the original trace was correctans or not.
+This gives a clearer picture than the overall accuracyans, which would be skewed by the proportion of traces that were originally correctans (15%) and incorrectans (85%).
+Scores represent the absolute differences in accuracyans. We perform backtracking on both correctans and incorrectans traces, as long as there is a mistake in one of the steps.
 accuracy
-refers to differences in accuracy on the set of traces whose original answer was correct.
+refers to differences in accuracyans on the set of traces whose original answer was correctans.
 Note that we take losses here because, despite the correct answer, there is a logical mistake in one of the steps.
@@
 accuracy
-is the same but for incorrect traces, so the answers may have been corrected, hence increasing accuracy.
-For example, for the word sorting task, 11.11% of traces that were originally correct became incorrect, while 23.53% of traces that were originally incorrect became correct.
-4.1.2 Discussion
-The scores show that the gains from correcting incorrect traces are larger than losses from changing originally correct answers.
+is the same but for incorrectans traces, so the answers may have been corrected, hence increasing accuracyans.
+For example, for the word sorting task, 11.11% of traces that were originally correctans became incorrectans, while 23.53% of traces that were originally incorrectans became correctans.
+4.2 Discussion
+The scores show that the gains from correcting incorrectans traces are larger than losses from changing originally correct answers.
 Additionally, while the random baseline also obtained improvements, they are considerably smaller than if the true mistake location was used.
 Note that tasks involving fewer steps are more likely to improve performance in the random baseline, as the true mistake location is more likely to be identified.
+Our results show that, with mistake location information available, LLMs can correct their own outputs and improve overall downstream performance.
+This suggests that the main bottleneck in self-correction methods is the identification of mistakes, rather than the correction process.
+This bottleneck can be overcome by using ground truth feedback (as in Reflexion (Shinn et al., 2023) or RCI (Kim et al., 2023)), or by training a classifier (see section 5).
 While our numbers do show that our gains are higher than our losses, it should be noted that changes in the overall accuracy depends on the original accuracy achieved on the task.
@@
 On the other hand, if the accuracy was 99%, the new accuracy would drop to 92.8%.
-As our dataset is highly skewed and only contains 45 correct traces per task, we leave to future work to assess the effectiveness of backtracking in a more comprehensive way.
-4.2 Backtracking with a simulated reward model
-We show in subsection 4.1 that backtracking can be used to correct CoT traces using gold mistake location labels.
-To explore what level of accuracy reward model is needed when gold labels are not available, we use backtracking with simulated reward models, designed to produce labels at different levels of accuracy.
-We use accuracy to refer to the accuracy of the simulated reward model at identifying mistake locations.
-For a given reward model at % accuracy, we use the mistake location from BIG-Bench Mistake % of the time.
+As our dataset is highly skewed and only contains 45 correctans traces per task, we leave to future work a more comprehensive assessment of backtracking, as well as the development of more sophisticated ways to incorporate mistake location information into the self-correction loop.
+5 Obtaining mistake location information with a trained classifier
+As shown in section 4, if mistake location information is available, LLMs can correct their own CoT traces and boost downstream performance.
+However, these experimental results are based on oracle labels, which are typically not available in downstream tasks.
+One possible solution is to obtain mistake location information from a smaller, trained classifier.
+If training data is available, one might ask why this approach is preferable to simply fine-tuning the larger, generator model.
+The reasons are:
+Training a small classifier is far more efficient in terms of computing resources and available data.
+Once the classifier is trained, it can be used with any LLM as the generator and be updated independently. This can be especially helpful with API-based LLMs that cannot be fine-tuned.
+The process of mistake finding is more interpretable than updating the weights of the generator model directly.
+It clearly pinpoints the location at which an error occurs, which can help the debugging process and allow faster development and iterations of models.
+In this section, we seek to answer two questions in the following subsections:
+5.1: What mistake-finding accuracy is required for backtracking to be effective?
+A trained classifier is unlikely to reach 100% mistake-finding accuracy.
+If backtracking is only effective when mistake location is 100% accurate, we would not be able to replace oracle labels with a trained classifier.
+5.2: Is it possible to improve on results in section 3 without in-domain training data?
+Sufficient in-domain training data typically guarantees a performance boost, but can be hard to obtain.
+We investigate whether mistake-finding in reasoning traces is transferable across tasks.
+If so, one can use BIG-Bench Mistake or similar datasets to fine-tune a mistake-finding classifier for other tasks.
+5.1 Minimum mistake finding accuracy
+To explore what level of mistake-finding accuracy is needed, we simulate classifiers at different levels of accuracy and run backtracking for each level.
+We use accuracymis to refer to the mistake-finding accuracy classifier, to differentiate from downstream task accuracyans.
+For a given classifier at % accuracyclf, we use the mistake location from BIG-Bench Mistake % of the time.
 For the remaining %, we sample a mistake location randomly.
@@
 We also ensure that the sampled location does not match the correct location.
-4.2.1 Results
 Results are shown in Figure 3.
 We can see that the losses in accuracy begins to plateau at 65%.
-In fact, for most tasks, accuracy is already larger than accuracy at around 60-70% accuracy.
+In fact, for most tasks, accuracy is already larger than accuracy at around 60-70% accuracyRM.
 This demonstrates that while higher accuracies produce better results, backtracking is still effective even without gold standard mistake location labels.
-4.3 Reward Modeling
-We perform a preliminary investigation of if mistake-finding can benefit from a dedicated reward model and if learning to find mistakes in a set of tasks can transfer to finding mistakes in out-of-distribution tasks.
-We fine-tuned a PaLM 2-XS-Otter model based on our available data for 20k steps and choose the checkpoint with the best validation results.
-We hold out one task for evaluation while training the reward model on the other 4 tasks.
-| Held-out task | Difference |
-| Word sorting | +11.66 |
-| Tracking shuffled objects | +19.67 |
-| Logical deduction | -0.67 |
-| Multi-step arithmetic | +4.00 |
-| Dyck languages | +22.59 |
-Note the reward model we train is significantly smaller than our inference model.
+5.2 Training a classifier on out-of-domain data
+We test whether mistake-finding can benefit from a dedicated classifier trained on out-of-distribution tasks.
+We fine-tune PaLM 2 Otter, a model much smaller than PaLM 2 Unicorn, with our BIG-Bench Mistake data for 20k steps and choose the checkpoint with the best validation results.
+For each task, we hold out the in-domain data for evaluation while training the classifier on the other 4 tasks.
+Held-out task | Trained classifier accuracymis (Otter) | 3-shot prompting accuracymis (Unicorn) | Difference |
+Word sorting | 22.33 | 11.67 | +11.66 |
+Tracking shuffled objects | 37.67 | 18.00 | +19.67 |
+Logical deduction | 6.00 | 6.67 | -0.67 |
+Multi-step arithmetic | 26.00 | 22.00 | +4.00 |
+Dyck languages | 33.57 | 10.98 | +22.59 |
 We show the relative improvements and losses in Table 7 vs.
-a zero-shot baseline on PaLM 2-L-Unicorn.
+a 3-shot baseline on PaLM 2 Unicorn (scores from section 3).
 We see gains for 4 out of 5 of the tasks.
-This provides initial indication that it maybe possible to train separate reward model classifiers to assist in backtracking and that these reward models do not have to be large.
-Further, a reward model can work on mistakes that are out-of-distribution.
-However, we believe more data may be necessary to improve results across the board on all tasks.
-We leave the collection of this larger dataset and a more rigorous investigation of the trade-offs of model size vs.
-performance of the reward model to future work.
-We also leave for future work the effect of backtracking iteratively with a reward model: for example, the generator model may make another mistake after backtracking for the first time, which can then be identified and corrected again.
+Note the classifiers we train are significantly smaller than our inference model, and is trained on out-of-domain data.
+This suggests that it may be possible to train classifiers to assist in backtracking, and that these classifiers do not have to be large.
+Further, such a classifier can work on out-of-distribution mistakes.
+Despite this, the performance of our classifiers do not meet the threshold required for effective backtracking, as demonstrated in subsection 5.1.
+We believe more data may be necessary to improve results across the board on all tasks.
+We leave to future work the collection of this larger dataset and a more rigorous investigation of the trade-offs of model size vs.
+performance of the classifier.
+We also leave for future investigation the effect of backtracking iteratively with a classifier: for example, the generator model may make another mistake after backtracking for the first time, which can then be identified and corrected again.
 6 Related work
-Datasets
+6.1 Datasets
 To our knowledge, the only publicly available dataset containing mistake annotations in LLM outputs is PRM800K (Lightman et al., 2023), which is a dataset of solutions to Olympiad-level math questions.
 Our dataset BIG-Bench Mistake covers a wider range of tasks to explore the reasoning capabilities of LLMs more thoroughly.
-Additionally, the generator LLM used in PRM800K has been fine-tuned on 1.5B math tokens as well as a dataset of step-by-step math solutions.
-For this paper, we wanted to explore few-shot in-context learning methods, which is typically used in real-world applications with API-based LLMs.
-Self-correction
+Additionally, our dataset explores the more general case of prompting API-based LLMs, whereas PRM800K uses a generator LLM that was heavily fine-tuned for math-specific problems.
+6.2 Self-correction
+The term self-correction can be applied to a wide variety of techniques.
 Pan et al.
-(2023) present a plethora of self-correction methods in recent literature.
-While their list includes training-time correction strategies such as RLHF (Ouyang et al., 2022) and self-improve (Huang et al., 2022), our backtracking method falls into the category of post-hoc correction, where the correction process is applied to outputs that have already been generated.
-Our paper focuses on correction of logical and reasoning errors, rather than stylistic or qualitative improvements.
-Previous post-hoc correction methods that are applied to reasoning errors include Reflexion (Shinn et al., 2023) and RCI (Kim et al., 2023), both of which cause performance deterioration when the oracle label is not used (Huang et al., 2023).
-Other methods such as Self-Refine (Madaan et al., 2023) and iterative refinement (Chen et al., 2023) focus on qualitative or stylistic improvements rather than correcting logical errors.
+(2023) present a survey of self-correction methods in recent literature.
+Broadly, proposed self-correction methods can vary in the following dimensions:
+Source of feedback
+Some techniques rely on external feedback inherent to the task such as code execution errors (Yasunaga and Liang, 2020; Chen et al., 2023a, 2024b), or feedback from humans (e.g.
+Chen et al., 2024a; Yuan et al., 2024).
+Some explicitly train a model to produce feedback (e.g.
+Madaan et al., 2021; Welleck et al., 2023; Bai et al., 2022b; Lee et al., 2023; Paul et al., 2023; Ouyang et al., 2022; Bai et al., 2022a; Ganguli et al., 2023), while others rely on prompting only (e.g.
+Shinn et al., 2023; Miao et al., 2024).
+It has been demonstrated that prompting-only setups can work well for stylistic or qualitative improvements (e.g.
+Madaan et al., 2023; Chen et al., 2023b), but would require external feedback for reasoning tasks (Huang et al., 2023; Shinn et al., 2023; Kim et al., 2023; Madaan et al., 2023).
+Time of correction
+Feedback can be incorporated at various points during the self-correction process.
+Some methods do so by updating weights during training time (e.g.
+Ouyang et al., 2022; Bai et al., 2022a; Ganguli et al., 2023); some do so during generation time (e.g.
+Weng et al., 2023; Dalvi Mishra et al., 2022; Xie et al., 2023); others apply correction to output that has already been generated (e.g.
+Saunders et al., 2022; Kim et al., 2023; Shinn et al., 2023).
+Our method falls into the final category (post-hoc correction), as it involves identifying an incorrect step in a complete CoT trace; however, it is also possible to apply a mistake-finding classifier at every step during generation.
 7 Conclusion
-In this paper, we describe and release our dataset BIG-Bench Mistake for mistake finding, and propose a backtracking method to correct logical errors in CoT style traces.
-We show that LLMs generally struggle with finding logical errors without external feedback, but argue that this feedback can come from a reward model instead.
-Finally, we demonstrate the effectiveness of backtracking, both with gold standard labels as well as with simulated reward models at lower levels of accuracy.
+In this paper, we investigate LLMs’ ability to find mistakes and correct outputs.
+We find that LLMs generally struggle to find mistakes, but, when given mistake location information, are able to correct outputs to boost performance.
+We therefore hypothesise that mistake finding is an important bottleneck preventing self-corrections strategies from performing well on reasoning tasks.
+We show initial evidence that a dedicated classifier for mistake finding can overcome this bottleneck.
+We train a small, baseline classifier on out-of-domain data and demonstrate improvement over few-shot prompting results.
+While the classifier does not reach the threshold required for effective backtracking, our results show that it is possible to improve on mistake-finding accuracy using standard machine learning techniques.
+We leave the development of more sophisticated methods to future work, and release our dataset BIG-Bench Mistake to encourage this direction of research.
 Limitations
@@
 We made this choice to minimise ambiguity and subjectivity during the mistake finding process, but further work needs to be done to determine the effectiveness of backtracking in a more realistic setting.
-Another limitation is that our paper does not evaluate backtracking on the original datasets on BIG-Bench, only showing results on the limited set that we sampled in a skewed manner, in order to maximise the value of the human annotators’ time.
-We leave the full evaluation to future work.
-Appendix A Implementational details
+Another limitation is that our paper does not experiment with backtracking on the original datasets on BIG-Bench, only showing results on the limited set that we sampled in a skewed manner, in order to maximise the value of the human annotators’ time.
+We leave the full evaluation to future work as this is beyond the scope of this paper, which is intended as a proof-of-concept to show the importance of mistake-finding.
+Acknowledgements
+We thank Vicky Zayats and Sian Gooding for their invaluable feedback, as well as members of the Google Research and Google DeepMind teams who gave continued support throughout the project.
+Appendix A Dataset details
+Our dataset, BIG-Bench Mistake, is available at https://github.com/WHGTyen/BIG-Bench-Mistake under the Apache License 2.0.
+The five tasks used in our dataset are based on BIG-Bench (Srivastava et al., 2022), also released under the Apache License 2.0.
+All five tasks are in the English language.
 A.1 3-shot CoT prompting to generate traces for BIG-Bench Mistake
-We use PaLM 2 L (Unicorn) to generate the traces used in BIG-Bench Mistake. All traces are generated at temperature .
+We use PaLM 2 (Unicorn) to generate the traces used in BIG-Bench Mistake. All traces are generated at temperature .
 Our prompts and examples can be found at https://github.com/WHGTyen/BIG-Bench-Mistake.
@@
 1.
-Example CoT traces in the prompt is broken up into smaller steps (typically one sentence per step). This is done so that mistake location information is more precise.
+Example CoT traces in the prompt are broken up into smaller steps (typically one sentence per step). This is done so that mistake location information is more precise.
 2.
@@
 It becomes difficult for annotators to determine whether a question mark is a mistake or not, because the correctness of the question mark is dependent on its interpretation.
-To minimise such ambiguity, the question mark notation is rewritten into text.
+To minimise such ambiguity, we replace the question mark notation with text descriptions of the objects.
 4.
@@
 Following Yao et al.
-(2022), we use the newline as the stop token, which generates one step with every generation call.
+(2022), we use the newline as the stop token, thereby generating one step with every generation call.
 We algorithmically append “Thought N:” before each step.
@@
 Appendix B Annotation
-We release our annotation guidelines at https://github.com/WHGTyen/BIG-Bench-Mistake.
+We release our annotation guidelines at https://github.com/WHGTyen/BIG-Bench-Mistake. Our annotators are recruited via our institution and contracted at the market rate in their country of residence.
 During annotation of the multistep arithmetic task, we found that the first CoT step given in the original BIG-Bench Hard prompt examples (Suzgun et al., 2022) was incorrect.
@@
 Figure 4 contains an example screenshot of the user interface.
-For every trace, we provide the input question as well as the target answer, with a note to be aware of errors that may occur in correct traces.
+For every trace, we provide the input question as well as the target answer, with a note to be aware of errors that may occur in correctans traces.
 Annotators can click on words to highlight the same word across the trace and the question text, which we found was particularly helpful for some tasks such as word sorting and tracking shuffled objects.
 Buttons on the right automatically become inactive if a previous step has been labelled as negative.
-| Model |
-Direct (trace) |
-Direct (step) |
-CoT (step) |
+Appendix C Training mistake-finding classifiers
+To train our mistake-finding classifiers (see subsection 5.2), we fine-tune PaLM 2 Otter on 4 of our 5 tasks, holding out one task for evaluation. This is done for each of our 5 tasks.
+All 5 models are fine-tuned for 20k steps with a batch size of 32.
+The learning rate is with a linear ramp and cosine decay.
+After 20k steps, we select the checkpoint with the best validation results.
+The number of steps trained for each model are shown in Table 8.
+Held-out task | Training steps |
+Word sorting | 6800 |
+Tracking shuffled objects | 8000 |
+Logical deduction | 9000 |
+Multi-step arithmetic | 10000 |
+Dyck languages | 10000 |
+All models are trained as a binary classifier on whether a CoT step is correct, given the task and previous steps.
+Due to the limited data, we include in training the CoT steps that occur after the first mistake step.
+These steps are considered incorrect for the purposes of training (despite not being human-annotated as such).
+Appendix D User interface
+Appendix E Benchmark scores
+Model | Direct (trace) | Direct (step) | CoT (step) |
 Word sorting | |||
-| GPT-4-Turbo |
-67.74 |
-38.24 |
-| GPT-4 |
-88.24 |
-82.35 |
-58.82 |
-| GPT-3.5-Turbo |
-100.00 |
-97.06 |
-20.59 |
+GPT-4-Turbo | 67.74 | 38.24 | – |
+GPT-4 | 88.24 | 82.35 | 58.82 |
+GPT-3.5-Turbo | 100.00 | 97.06 | 20.59 |
+Gemini Pro | 44.12 | – | – |
+PaLM 2 Unicorn | 100.00 | 73.53 | 35.29 |
 Tracking shuffled objects | |||
-| GPT-4-Turbo |
-90.00 |
-77.50 |
-| GPT-4 |
-82.50 |
-82.50 |
-80.00 |
-| GPT-3.5-Turbo |
-67.50 |
-0.00 |
-0.00 |
+GPT-4-Turbo | 90.00 | 77.50 | – |
+GPT-4 | 82.50 | 82.50 | 80.00 |
+GPT-3.5-Turbo | 67.50 | 0.00 | 0.00 |
+Gemini Pro | 12.50 | – | – |
+PaLM 2 Unicorn | 100.00 | 85.00 | 47.50 |
 Logical deduction | |||
-| GPT-4-Turbo |
-100.00 |
-83.33 |
-| GPT-4 |
-100.00 |
-100.00 |
-0.00 |
-| GPT-3.5-Turbo |
-100.00 |
-50.00 |
-100.00 |
+GPT-4-Turbo | 100.00 | 83.33 | – |
+GPT-4 | 100.00 | 100.00 | 0.00 |
+GPT-3.5-Turbo | 100.00 | 50.00 | 100.00 |
+Gemini Pro | 33.33 | – | – |
+PaLM 2 Unicorn | 100.00 | 100.00 | 50.00 |
 Multistep arithmetic | |||
-| GPT-4-Turbo |
-57.69 |
-40.32 |
-| GPT-4 |
-53.23 |
-46.77 |
-27.42 |
-| GPT-3.5-Turbo |
-96.77 |
-79.03 |
-58.06 |
+GPT-4-Turbo | 57.69 | 40.32 | – |
+GPT-4 | 53.23 | 46.77 | 27.42 |
+GPT-3.5-Turbo | 96.77 | 79.03 | 58.06 |
+Gemini Pro | 83.87 | – | – |
+PaLM 2 Unicorn | 83.87 | 93.55 | 29.03 |
 Dyck languages | |||
-| GPT-4-Turbo |
-96.42 |
-30.00 |
-| GPT-4 |
-98.41 |
-78.57 |
-13.79 |
-| GPT-3.5-Turbo |
-95.74 |
-4.76 |
-0.00 |
-| Model |
-Direct (trace) |
-Direct (step) |
-CoT (step) |
+GPT-4-Turbo | 96.42 | 30.00 | – |
+GPT-4 | 98.41 | 78.57 | 13.79 |
+GPT-3.5-Turbo | 95.74 | 4.76 | 0.00 |
+Gemini Pro | 0.00 | – | – |
+PaLM 2 Unicorn | 100.00 | 80.95 | 19.05 |
+Model | Direct (trace) | Direct (step) | CoT (step) |
 Word sorting | |||
-| GPT-4-Turbo |
-32.71 |
-32.33 |
-| GPT-4 |
-28.20 |
-39.47 |
-30.83 |
-| GPT-3.5-Turbo |
-0.00 |
-4.51 |
-15.04 |
+GPT-4-Turbo | 32.71 | 32.33 | – |
+GPT-4 | 28.20 | 39.47 | 30.83 |
+GPT-3.5-Turbo | 0.00 | 4.51 | 15.04 |
+Gemini Pro | 6.39 | – | – |
+PaLM 2 Unicorn | 0.38 | 9.02 | 11.28 |
 Tracking shuffled objects | |||
-| GPT-4-Turbo |
-31.54 |
-59.23 |
-| GPT-4 |
-59.14 |
-62.69 |
-92.31 |
-| GPT-3.5-Turbo |
-1.17 |
-1.92 |
-21.92 |
+GPT-4-Turbo | 31.54 | 59.23 | – |
+GPT-4 | 59.14 | 62.69 | 92.31 |
+GPT-3.5-Turbo | 1.17 | 1.92 | 21.92 |
+Gemini Pro | 41.54 | – | – |
+PaLM 2 Unicorn | 5.38 | 19.23 | 56.92 |
 Logical deduction | |||
-| GPT-4-Turbo |
-20.81 |
-74.83 |
-| GPT-4 |
-39.46 |
-67.01 |
-10.54 |
-| GPT-3.5-Turbo |
-0.00 |
-24.83 |
-7.82 |
+GPT-4-Turbo | 20.81 | 74.83 | – |
+GPT-4 | 39.46 | 67.01 | 10.54 |
+GPT-3.5-Turbo | 0.00 | 24.83 | 7.82 |
+Gemini Pro | 8.16 | – | – |
+PaLM 2 Unicorn | 4.76 | 36.73 | 11.22 |
 Multistep arithmetic | |||
-| GPT-4-Turbo |
-34.27 |
-44.12 |
-| GPT-4 |
-41.60 |
-41.60 |
-44.54 |
-| GPT-3.5-Turbo |
-0.00 |
-12.18 |
-16.81 |
+GPT-4-Turbo | 34.27 | 44.12 | – |
+GPT-4 | 41.60 | 41.60 | 44.54 |
+GPT-3.5-Turbo | 0.00 | 12.18 | 16.81 |
+Gemini Pro | 5.46 | – | – |
+PaLM 2 Unicorn | 5.88 | 2.94 | 22.27 |
 Dyck languages | |||
-| GPT-4-Turbo |
-6.99 |
-28.46 |
-| GPT-4 |
-7.37 |
-40.81 |
-43.91 |
-| GPT-3.5-Turbo |
-1.28 |
-6.05 |
-2.08 |
+GPT-4-Turbo | 6.99 | 28.46 | – |
+GPT-4 | 7.37 | 40.81 | 43.91 |
+GPT-3.5-Turbo | 1.28 | 6.05 | 2.08 |
+Gemini Pro | 2.25 | – | – |
+PaLM 2 Unicorn | 0.38 | 6.43 | 17.77 |